OpenStudy (anonymous):

Find the solution of this system of equations. Separate the x- and y- values with a comma. x=-5+y 11x-y=-25

OpenStudy (anonymous):

Best answer gets a medal

OpenStudy (anonymous):

@Chiko_1278

OpenStudy (anonymous):

y=x+5 11x-x-5=-25 10 x=-25+5 10 x=-20 x=-20/10=-2 y=?

OpenStudy (anonymous):

5?